(2x/5) + 3 = (7x/10)
Answer by checkley79(3341)

(2x/5) + 3 = (7x/10)
2X/5-(7X/10=-3
(2*2X/5-7X)/10=-3
(4X-7X)/10=-3
-3X/10=3
-3X=-30
X=-30/-3
X=30/3
X=10 ANS.
PROOF:
(2*10)/5+3=(7*10)/10
20/5+3=70/10
4+3=7
7=7